OpenStudy (anonymous):

Age Ticket Price Under three years free Three to twelve years $36 Over twelve years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged twelve and under, of which some children were under three years. If x represents the number of children undr three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (92 − x)36 + (92 − 40) ∙ 48 = C (40 + x)36 + (92 − (40 + x)) ∙ 48 = C

OpenStudy (anonymous):

(92 + x)36 + (92 − (40 + x)) ∙ 48 = C (40 − x)36 + (92 − 40) ∙ 48 = C

OpenStudy (anonymous):

(92 − x)36 + (92 − 40) ∙ 48 = C (40 + x)36 + (92 − (40 + x)) ∙ 48 = C